6. Select the set of equations that represents the following situation: The sum of two numbers is 77, and their difference is nineteen less than the smaller number. What are the two numbers?
which one? A,B,C, or D?
(A) 2y = 77; 77 = x – 19
(B) 2x = 77; 77 = x + 19
(C) x + y = 77; x – y = 19 – x
(D) x + y = 77; y – x = x – 19

Answer by deborabr(181)

(D)x + y = 77; y – x = x – 19
because
The sum of two numbers is 77
so x+y=77
heir difference is nineteen less than the smaller number
the difference is the bigger number less the smaller number and if we suppose that y is the greater number we have: y-x
and nineteen less than the smaller number means: x-19